a concave lens is in contact with convex lens.The magnitude of the ratio of their powers is 2/3.Their equivalent focal length is 30cm.What are their individual focal lengths(in cm).

a)-75cm,+50cm
b)-15cm,10cm
c)+75cm,-50cm
d)75cm,50cm
 
Grade:

1 Answers

Eshan
5 years ago
Dear student,

P=1/f
\dpi{80} P_1/P_2=-2/3
\implies f_1=-\dfrac{3}{2}f_2
AlsoP=P_1+P_2
\dfrac{1}{30}=\dfrac{1}{f_1}+\dfrac{1}{f_2}=-\dfrac{2}{3f_2}+\dfrac{1}{f_2}
Hence findf_1,f_2